OpenStudy (anonymous):

which is equal to 6^-8 / 6^-2

OpenStudy (cwrw238):

when dividing this you subtract the powers x^a / x^b = x^(a-b)

OpenStudy (anonymous):

6^-6

OpenStudy (cwrw238):

so its 6^(-8 - (-2))

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

can u help me with one more?

OpenStudy (cwrw238):

you - i have got 5 minutes

OpenStudy (cwrw238):

yea I meant!!

OpenStudy (cwrw238):

use the following rule for indices (x^a)^b = x^ab

OpenStudy (anonymous):

12^15 ?

OpenStudy (anonymous):

@cwrw238